OpenStudy (anonymous):

A farmer hears a weather report saying that temperatures will be decreasing approximately 5 degrees every hour for the next 4 hours. What signed number represents the total change in temperature expected over the next 4 hours?

OpenStudy (anonymous):

if it decreases how 5 degrees every hour for 4 hours you'll get 1-hour=5 2-hours=5+5 3-hours=5+5+5 4-hours=....